solve for x and y
5x+3y=11
-9x-4y=-17

Answer by stanbon(75887)

1st: 5x+3y=11
2nd: -9x-4y=-17
---------
Multiply the 1st by 4 and the 2nd by 3 to get:
3rd: 20x + 12y = 44
4th:-27x - 12y = -51
----------------
Add 3rd and 4th to solve for x:
-7x = -7
x = 1
------------
Substitute into 1st (or 2nd) to solve for y:
5 + 3y = 11
3y = 6
y =2
--------------
Check this answer in 2nd.
Cheers,
Stan H.